OpenStudy (anonymous):

Find lim 1-sec^2x over cosx-1, x->0.

OpenStudy (anonymous):

maybe some algebra first, since this is like \[\frac{1-\frac{1}{a^2}}{a-1}=\frac{a+1}{x^2}\] so you are starting with \[\frac{\cos(x)+1}{\cos^2(x)}\]

OpenStudy (anonymous):

replace x by 0 and you will get the answer

OpenStudy (anonymous):

i get 2

OpenStudy (nenadmatematika):